SUB-DOPPLER SPECTROSCOPY OF H_{3}^{+}

\underline{JAMES N. HODGES}, ADAM J. PERRY, BRIAN M. SILLER, Department of Chemistry, University of Illinois, Urbana, IL 61801; BENJAMIN J. MCCALL, Departments of Chemistry, Astronomy, and Physics, University of Illinois, Urbana, IL 61801.

Spectroscopy of H_{3}^{+} is of fundamental interest for advancing ab initio efforts to calculate spectra with high precision and accuracy. H_{3}^{+} is the simplest polyatomic ion, which is why it is an excellent benchmark for theory. In order to perform calculations with spectroscopic accuracy, relativistic and non-adiabatic corrections to the Born-Oppenhiemer approximation must be included; calculations with these considerations agree to within hundredths of a wavenumber.^{a} Increasing the precision of the calculations further will require a treatment of quantum electrodynamic effects, as has already been implemented for the diatomic case,^{b} and testing these calculations will require higher-precision experimental data to guide ab initio calculations.

Noise Immune Cavity Enhanced Optical Heterodyne Velocity Modulation Spectroscopy, or NICE-OHVMS^{c}^{d}, is a highly sensitive, highly precise technique that we have employed to observe transitions in the ν_{2} fundamental band of H_{3}^{+}. It combines the advantages of cavity enhancement and heterodyne detection with the ion-neutral discrimination afforded by velocity modulation. Combining a cavity with a high power mid-infrared light source, we can saturate rovibrational transitions. The resulting Lamb dips may be fit in order to determine line centers to a much higher precision than is possible for ordinary Doppler broadened profiles. Additionally, a frequency comb is used to surpass the limited accuracy and precision of a wavemeter. Here we present the results from comb calibrated H_{3}^{+} transitions observed via NICE-OHVMS. Precision and accuracy of ∼ 1 MHz were achieved representing the most accurate and precise H_{3}^{+} line list that has been obtained to date.

aO. L. Polyansky, J. Tennyson, J. Chem. Phys. (1999), 110, 5056–5064.

bJ. Komasa, et al. J. Chem. Theor. Comp. (2011), 7, 3105–3115.

cB. M. Siller, et al. Opt. Express (2011), 19, 24822–7.

dK. N. Crabtree, et al. Chem. Phys. Lett. (2012), 551, 1–6.